find λ if the distance between (λ,1)and (1,-λ) is 2root5
 

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Nikhi
 
using distance formula
 
\large \sqrt(\large \lambda –1)2 + (1+ \large \lambda)2 = 2 \large \sqrt5
 
\large \sqrt(2 \large \lambda2 + 2) = 2 \large \sqrt5
on squaring both the sides
(2 \large \lambda2 + 2) = 20
 \large \lambda2 = 18/2 = 9
 \large \lambda = + 3 and  – 3
 
Regards
Arun (askIITians forum expert)